] A quantum particle is represented by a normalized wave function f(x) = √15 (a? _ x²) is the 4 a 512 region - a< x <a & val=Explain in detail for thumbs up. Quantum mechanics question
Answer #1

2 49/ - X IS V () = VTS (a-x) <P) = S (au) (-i 13.) (a*x) da Toas (ax) (-2x) da - Isch + 9 2 I . <pa) (a47) (-imej (a+e) daOp = √ <p²) - (p) 2 so uncertainty in P is 5